Grade 12th passAlgebra

Find the cost of polishing a circular table-top of diameter 1.6 m, if the rate of polishing is ₹15/m2 . (Take π = 3.14)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that,
Diameter of the circular table-top = 1.6 m
We know that, radius (r) = d/2
= 1.6/2 = 0.8 m
Then,
Area of the circular table-top =πr2
= 3.14 × 0.82
= 3.14 × 0.8 ×0.8
= 2.0096 m2
Cost for polishing 1 m2area =₹15
Cost for polishing 2.0096 m2 area =₹15 × 2.0096 [given]
=₹30.144
Hence, the Cost for polishing 2.0096 m2 area is₹30.144.

Thanks